ganeshie8 (ganeshie8):

Is it possible to factor an mxn matrix into a product of two matrices mxd and dxn, for every d ?

ganeshie8 (ganeshie8):

When \(d=1\) and \(A = \begin{bmatrix} 1&2\\2&4 \end{bmatrix}\) : \[\begin{bmatrix} 1&2\\2&4 \end{bmatrix} = \begin{bmatrix} 1\\2 \end{bmatrix} * \begin{bmatrix} 1&2 \end{bmatrix}\]

ganeshie8 (ganeshie8):

When \(d=2\) and \(A = \begin{bmatrix} 1&2\\2&4 \end{bmatrix}\) : \[\begin{bmatrix} 1&2\\2&4 \end{bmatrix}=\begin{bmatrix}1&0\\1&1 \end{bmatrix} * \begin{bmatrix} 1&2 \\1&2\end{bmatrix}\]

ganeshie8 (ganeshie8):

When \(d=3\) and \(A = \begin{bmatrix} 1&2\\2&4 \end{bmatrix}\) : \[\begin{bmatrix} 1&2\\2&4 \end{bmatrix}=\begin{bmatrix}1&0&0\\2&-1&-1 \end{bmatrix} * \begin{bmatrix} 1&2 \\1&2\\1&2\end{bmatrix}\]

ganeshie8 (ganeshie8):

looks i can keep factoring like that for every \(d\) ?

myininaya (myininaya):

i think you can too... i was looking at the row vectors of your second matrix in your factor form it is very repetitive [1 2] I think this is because the matrix A column vectors are linear dependent I think I remember the right word there like column 1 times 2 is column 2 in matrix A I'm just assuming it wouldn't be as pretty if the columns weren't linear dependent

myininaya (myininaya):

oh yeah the rows of matrix A are linear dependent too

myininaya (myininaya):

how did you factor it so fast

ganeshie8 (ganeshie8):

Yeah, I took the rows to be dependent so that factoring will be easy

ganeshie8 (ganeshie8):

fixed a typo When \(d=3\) and \(A = \begin{bmatrix} 1&2\\2&4 \end{bmatrix}\) : \[\begin{bmatrix} 1&2\\2&4 \end{bmatrix}=\begin{bmatrix}1&0&0\\2&-1&\color{red}{1} \end{bmatrix} * \begin{bmatrix} 1&2 \\1&2\\1&2\end{bmatrix}\]

ganeshie8 (ganeshie8):

I was looking at the row picture while factoring : first row of AB is obtained by taking linear combinations of the rows of B with the weights set by the first row of A

ganeshie8 (ganeshie8):

\[\begin{bmatrix}1&0&0\\2&-1&1 \end{bmatrix} * \begin{bmatrix} 1&2 \\1&2\\1&2\end{bmatrix}\] To obtain the "second" row of the above product, we take the combinations of rows of the second matrix like this : 2(first row) - 1(second row) + 1(third row)

ganeshie8 (ganeshie8):

\[\large 2 [1~~~2] - 1[1~~~2] +1[1~~~2] \]

ganeshie8 (ganeshie8):

\[\large [2~~~4] - [1~~~2] +[1~~~2] \] \[\large [2~~~4] \] That must be the second row of the product

myininaya (myininaya):

well i know how to do matrix multiplication i thought you were just using same fast method to factor matrix A i guess you were just working backwards

ganeshie8 (ganeshie8):

No, I don't have any particular method to factor. Just guessing.. I find multuplying like this way easier than working it element by element

ganeshie8 (ganeshie8):

For no reason I feel factoring like this should not be possible for every \(d\)

ganeshie8 (ganeshie8):

Let me try an invertible matrix may be

ganeshie8 (ganeshie8):

When \(d=2\) and \(A = \begin{bmatrix} 1&2\\2&1 \end{bmatrix}\) : \[\begin{bmatrix} 1&2\\2&5 \end{bmatrix} = \begin{bmatrix} 1&0\\2 &1\end{bmatrix} * \begin{bmatrix} 1&2\\0&1 \end{bmatrix}\]

ganeshie8 (ganeshie8):

Definitely, we cannot factor a 2x2 invertible matrix into a product of 2x1 and 1x2 matrices

ganeshie8 (ganeshie8):

It seems the minimum value of \(d\) must be rank of the matrix

ganeshie8 (ganeshie8):

When \(d=3\) and \(A = \begin{bmatrix} 1&2\\2&5 \end{bmatrix}\) : \[\begin{bmatrix} 1&2\\2&5 \end{bmatrix} = \begin{bmatrix} 1&0&0\\2 &1&0\end{bmatrix} * \begin{bmatrix} 1&2\\0&1 \\\spadesuit & \clubsuit \end{bmatrix}\]

ganeshie8 (ganeshie8):

When \(d=4\) and \(A = \begin{bmatrix} 1&2\\2&5 \end{bmatrix}\) : \[\begin{bmatrix} 1&2\\2&5 \end{bmatrix} = \begin{bmatrix} 1&0&1&1 \\2 &1&1&1\end{bmatrix} * \begin{bmatrix} 1&2\\0&1 \\\spadesuit & \clubsuit \\ -\spadesuit & -\clubsuit \end{bmatrix}\]

ganeshie8 (ganeshie8):

guess I'm looking for a proof of this : when \(d\ge rank(A)\), there exist matrices \(X, Y\) such that \[\large A_{m\times n} = X_{m\times d}*Y_{d\times n}\]

OpenStudy (christos):

I noticed that when A is singular and d = 1 then the vectors of the dot product are always in the row space of A ... but the row space of A is only 1 dimension .. so I guess the logic behind it is that you can somehow exploit the fact that the second row of A give no additional information ..

ganeshie8 (ganeshie8):

That's a nice observation. Basically, you're saying that the rows of XY always lie in the row space of Y.

ganeshie8 (ganeshie8):

Also, columns of XY always lie in the column space of X.

OpenStudy (christos):

Yeah . Also when d=1 there is no vector a that if you dot it with a vector b is going to give you an invertible matrix, so if A is invertible to begin with , then a*b can never be A

OpenStudy (christos):

cause if d=1 a/b will always have dependent collumns/rows

ganeshie8 (ganeshie8):

Ahh nice nice Is it right to say that the rank of XY must be less than the ranks of X and Y ?

OpenStudy (christos):

I guess yea

OpenStudy (christos):

would be interesting to have a formal proof , but this is more challenging :D

OpenStudy (kainui):

I think it helps if you think of these as maps and whether or not they can be factored or not. \[A_{mn} =B_{md}C_{dn}\] since A maps column vectors from \(\mathbb{R}^n\) to \(\mathbb{R}^m\) (and m to n for row vectors if left multiplied) then in order to not lose information, \(d \ge \min(m,n)\) I think.

ganeshie8 (ganeshie8):

Nice, assuming either m or n is the rank of A. if the rank itself is less than m or n, d can be made less right ?

OpenStudy (kainui):

Yeah I guess I was assuming the smaller of m or n was the rank of A, haha whoops